a boy is three years younger than his sister. if his age three years ago was two thirds her age at that time. what are their present ages ?
Answer by ptaylor(2198)

Let x=sister's present age; x-3 is sister's age three years ago
Then x-3=boy's present; x-6=boy's age three years ago
Now we are told the following
x-6=(2/3)(x-3) multiply each term by 3
3x-18=2x-6 add 18 to and subtract 2x from each side
3x-2x=12
x=12---------------------sister's present
x-3=12-3=9-----------------boy's present age
CK
9-3=(2/3)(12-3)
6=(2/3)(9)=6
6=6
Does this help???----ptaylor